### v1

@@ Abstract
information for the production of the at collider
@@
A series of exotic hadron candidates, which can not be accommodated by the potential model Eichten et al. (1978, 1980); Godfrey and Isgur (1985), were observed experimentally in the charmonium energy region during the past decades.
@@
of molecule, quark-gluon hybrid, or tetraquark states Olsen et al. (2018). The well-known state was firstly observed
@@
as its mass is quite near mass threshold.
@@
and the might be a good candidate for due to the similar mass as
that from potential model prediction Eichten et al. (1978). The other interpretations such as a tetraquark candidate Maiani et al. (2005); Esposito et al. (2015) is also possible.
However, these is still no solid conclusion for the nature of .
@@
What intriguing is that the might originate from the radiative transition of an excited vector state , which for the first time brings together two
charmonium-like states and hints commonality for their underlying nature Olsen et al. (2018). Recently, the
process is observed at BESIII Ablikim et al. (2023a), and the production cross section shows potential enhancement near 4.75 GeV.
@@
of production will be achieved and further to deeply explore its nature Maiani et al. (2020).
@@
The meson is reconstructed via decay, while the via decay
@@
processes, the ISR production of vector charmonium(-like) states,
and the continuum processes incorporated in kkmc. All particle decays are modelled with evtgen Lange (2001); Ping (2008) using branching fractions
@@
the remaining unknown charmonium decays are modelled with lundcharm Chen et al. (2000); Yang et al. (2014).
@@
The energy deposited in the EMC is utilized to separate electrons from muons. For both
@@
from kaons, particle identification (PID), which combines measurements of the specific ionization
@@
likelihoods for each hadron hypothesis is used.
@@
by maximizing the figure-of-merit (FOM) ,
@@
The signal window of resonance is set as
GeV (mass resolution is 7 MeV) with barely efficiency loss according to signal MC events.
@@
same as wide as the signal region.
After applying all the selection criteria, there is no
background in the 6-track events, as indicated by inclusive MC sample.
Figure 1 shows the invariant mass of lepton pairs and kaon pairs
distributions for 6 track events from the full data set.
@@
signal MC sample, and the blue filled histograms are inclusive MC sample.
@@
A one-constraint (1C) kinematic fit is performed to the 5-track events,
by constraining the mass of missing particle to the nominal mass of the
@@
which is optimized by maximizing FOM=.
@@
of resonance is set to be GeV (mass resolution is 8 MeV) and
@@
to identify muons. At least one of muon candidate should have a hit depth 30 cm in the MUC.
@@
and are the four-momenta of the initial colliding beams and the system.
@@
Dots with error bars are 4.914 and 4.946 GeV data, the red histograms are signal MC sample, and
the blue filled histograms are inclusive MC sample.
@@
the distribution from 6-track events and 5-track events at each c.m. energy. The invariant mass = is defined, which
partly helps cancel the resolution effect of lepton pairs.
@@
no dominant background survives for both 6-track events and 5-tracks events.
The study of the and mass sideband events also shows the background level is low and we neglect the
@@
Dots with error bars are data, the red histograms are signal MC sample,
the blue filled histograms are inclusive MC sample and
@@
Due to no event observed in the signal region, we report an upper limit on the Born cross section by combining the 5-track events and 6-track events using a counting method.
The number of signal events is determined by counting the
number of events in the signal region from 5-track events and 6-track events, which is defined as GeV.
@@
and GeV from 5-track events and 6-track events is subtracted
(at the sideband region is defined as GeV due to the limitation of kinematics).
The upper limit on the number of signal yield at the 90 confidence level (C.L.) is obtained
using a frequentist method with an unbounded profile likelihood treatment Rolke et al. (2005).
@@
the efficiency (sum of 5-track events and 6-track events) follows a Gaussian
@@
Table 1: The upper limit on the product of Born cross section
@@
region, is the number of observed events in sideband region,
is the upper limit on the number of observed signal at the 90 C.L.,
@@
The uncertainty from luminosity measurement is estimated to be less than 0.66 using
@@
The uncertainties of decay branching fractions are quoted from the PDG Workman et al. (2022).
The uncertainty of tracking efficiency for high momentum leptons is assigned to be 1 per track according
@@
In this measurement, both one kaon events
@@
The total uncertainty of tracking efficiency
of the kaon is 1.0 for 5-track events and 2.0 for 6-track events.
@@
the kaon PID uncertainty is 1.0 for 5-track events and 2.0 for 6-track events, too.
@@
The uncertainty of MUC hit depth is studied using control sample of Ablikim et al. (2023b). The
@@
the requirement of hit depth in the MUC is taken as the systematic uncertainty.
@@
caused by kinematic fit. The difference in efficiency with and
@@
The systematic uncertainty of radiative correction factor is studied
@@
is modelled by a 1 cos
distribution, the efficiency difference with respect to PHSP is taken as the systematic uncertainty.
@@
In the measurement, two sub-data samples, i.e. the 5-track events and 6-track events
are reconstructed. The same source of systematic uncertainty
contribute differently to two sub-data samples.
To combine the systematic uncertainty, we take the weighted average of the systematic
uncertainties in two sub-data samples, which follows:
@@
and are the weight and systematic uncertainty for the th sub-data sample, is the
efficiency for the th sub-data sample. For the soft kaon decay events (),
its weight contributes to 5-track sub-data sample for mass window, MUC hit depth requirement, kinematic fit;
otherwise its weight contributes to 6-track sub-data sample.
@@
Table 2 and Table 3 summarize all the
@@ 6 Tracks
upper limits at the 90% C.L. on the product of the Born cross section
@@
Ablikim et al. (2023a, ).
These measurements provide important inputs to the production of at collider, and help constrain the possible component in the
@@
With the upgrade of the BEPCII Ablikim et al. (2020b) project, more data in this energy region is expected and a more comprehensive study of the production will be achieved, which finally help reveal the nature of .
@@ Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2020YFA0406300, 2020YFA0406400, 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11835012, 11935015, 11935016, 11935018, 11961141012, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; CAS Key Research Program of Frontier Sciences under Contracts Nos. QYZDJ-SSW-SLH003, QYZDJ-SSW-SLH040; 100 Talents Program of CAS;
Project
No. ZR2022JQ02 supported by Shandong Provincial
@@
The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; European Union’s Horizon 2020 research and innovation programme under Marie Sklodowska-Curie grant agreement under Contract No. 894790; German Research Foundation DFG under Contracts Nos. 455635585, Collaborative Research Center CRC 1044, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
Chatrchyan et al. (2013)S. Chatrchyan et al. (CMS Collaboration), “Measurement of the
(3872) Production Cross Section Via Decays to in
collisions at = 7 TeV,” JHEP 04, 154 (2013), arXiv:1302.3968
[hep-ex] .
Cao et al. (2020)P. Cao et al., “Design and construction of the new BESIII endcap Time-of-Flight system with
MRPC Technology,” Nucl. Instrum. Meth. A 953, 163053 (2020).
Chen et al. (2000)J. C. Chen, G. S. Huang,
X. R. Qi, D. H. Zhang, and Y. S. Zhu, “Event generator for and
decay,” Phys. Rev. D 62, 034003 (2000).
Yang et al. (2014)R. L. Yang, R. G. Ping, and H. Chen, “Tuning and Validation of the Lundcharm
Model with Decays,” Chin. Phys. Lett. 31, 061301 (2014).
Actis et al. (2010)S. Actis et al. (Working Group on Radiative
Corrections, Monte Carlo Generators for Low Energies), “Quest for precision in hadronic cross sections
at low energy: Monte Carlo tools vs. experimental data,” Eur. Phys. J. C 66, 585–686 (2010), arXiv:0912.0749 [hep-ph] .

### v2

@@ Abstract
information for the production of the at colliders
@@
The quark model categorizes hadrons into two types: mesons with a quark and an antiquark, and baryons with three quarks.
Within the framework of the theory of strong interactions, quantum chromodynamics (QCD),
it also allows the existence of more complex structures, generically called exotic hadrons.
A series of exotic hadron candidates, which cannot be accommodated by the potential model Eichten et al. (1978, 1980); Godfrey and Isgur (1985),
were observed experimentally in the charmonium energy region during the past decades.
@@
of molecule, quark-gluon hybrid, or tetraquark states Olsen et al. (2018). The well-known state was first observed
@@
since its mass is quite near mass threshold.
@@
and the might be a good candidate for since its mass is similar to the
potential model prediction Eichten et al. (1978). Other interpretations such as a tetraquark candidate Maiani et al. (2005); Esposito et al. (2015) is also possible.
However, there is still no solid conclusion for the nature of .
@@
What is intriguing is that the might originate from the radiative transition of an excited vector state , which for the first time brings together two
charmoniumlike states and hints commonality for their underlying nature Olsen et al. (2018). Recently, the
process was observed at BESIII Ablikim et al. (2023a), and the production cross section shows potential enhancement near 4.75 GeV.
@@
of production will be achieved Maiani et al. (2020).
@@
The meson is reconstructed via decays, while the is found using decays
@@
processes, the ISR production of vector charmonium(like) states,
and the continuum processes incorporated in kkmc. All particle decays are modeled with evtgen Lange (2001); Ping (2008) using branching fractions
@@
the remaining unknown charmonium decays are modeled with lundcharm Chen et al. (2000); Yang et al. (2014).
@@
The energy deposited in the EMC is used to separate electrons from muons. For both
@@
from kaons, particle identification (PID) combines measurements of the specific ionization
@@
likelihoods for each hadron hypothesis.
@@
by maximizing the figure-of-merit
@@
The signal window of the resonance is set as
GeV (mass resolution is 7 MeV) with barely any efficiency loss according to signal MC events.
@@
as wide as the signal region.
After applying all of the selection criteria, there is no
background in the 6-track events, as indicated by the inclusive MC sample.
Figure 1 shows the invariant mass distributions of the lepton and kaon pairs
for 6-track events from the full dataset.
@@
the signal MC sample, and the blue filled histograms are the inclusive MC sample.
@@
A one-constraint (1C) kinematic fit is performed to these 5-track events,
by constraining the mass of the missing particle to the nominal mass of the
@@
which is optimized by maximizing the FOM defined in Eq. (Search for the process at BESIII).
@@
of the resonance is set to be GeV (mass resolution is 8 MeV) and
@@
to identify muons. At least one of the muon candidates should have a hit depth 30 cm in the MUC.
@@
and denoted the four-momenta of the initial colliding beams and the system.
@@
Dots with error bars are 4.914 and 4.946 GeV data, the red histograms are the signal MC sample, and
the blue filled histograms are the inclusive MC sample.
@@
the distribution from 6- and 5-track events at each c.m. energy. The invariant mass = is defined, which
partly helps to cancel the resolution effect of the lepton pairs.
@@
no dominant background survives for both 6- and 5-tracks events.
The study of the and mass sideband events also shows the background level is low and so we neglect the
@@
Dots with error bars are data, the red histograms are the signal MC sample,
the blue filled histograms are the inclusive MC sample and
@@
Due to no events observed in the signal region, we report an upper limit on the Born cross section
at the 90 confidence level (CL) using a frequentist method with an unbounded profile likelihood treatment Rolke et al. (2005).
The number of signal events is determined by the signal region for 6- and 5-track events,
which is defined as GeV.
@@
and GeV from 6- and 5-track events is subtracted
(at the sideband region is defined as GeV
due to the limitation of kinematics).
@@
the efficiency (sum of 6- and 5-track events) follows a Gaussian
@@
Table 1: The upper limit on the product of the Born cross section
@@
region, is the number of observed events in the sideband region,
is the upper limit on the number of observed signal at the 90 CL,
@@
The uncertainty from the luminosity measurement is estimated to be less than 0.66 using
@@
The uncertainties of the decay branching fractions are quoted from the PDG Workman et al. (2022).
The uncertainty of the tracking efficiency for high momentum leptons is assigned to be 1 per track according
@@
In this measurement, both one-
@@
The total uncertainty of the kaon tracking efficiency
is 1.0(2.0) for 5-track(6-track) events.
@@
the kaon PID uncertainty is 1.0(2.0) for 5-track(6-track) events, too.
@@
The uncertainty of the MUC hit depth is studied using the control sample of Ablikim et al. (2023b). The
@@
the requirement of the hit depth in the MUC is taken as the systematic uncertainty.
@@
caused by the kinematic fit. The difference in efficiency with and
@@
The systematic uncertainty of the radiative correction factor is studied
@@
is modeled by a 1 cos
distribution, where the efficiency difference with respect to PHSP is taken as the systematic uncertainty.
@@
In the measurement, two subdata samples, i.e. the 6- and 5-track events
are reconstructed. The same source of systematic uncertainties contribute to the two
subdata samples, namely the 5- and 6-track events. These are combined
by taking the weighted average of their systematic uncertainties: Eq. (Search for the process at BESIII) etc.
@@
while , and are the weight, efficiency and systematic uncertainty
for the th subdata sample, respectively. For the soft kaon decay events (),
its weight contributes to the 5-track subdata sample for the mass window, the MUC hit depth requirement and the kinematic fit;
otherwise its weight contributes to the 6-track subdata sample.
@@
Table 2 and Table 3 summarize all of the
@@
upper limits at the 90% CL on the product of the Born cross section
@@
Ablikim et al. (2023a, 2024).
These measurements provide important inputs to the production of at colliders, and help constrain the possible component in the
@@
With the upgrade of the BEPCII Ablikim et al. (2020b) project, more data in this energy region is expected and a more comprehensive study
of the production will be achieved, which will hopefully reveal the nature of state.
@@ Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts No. 2020YFA0406300, No. 2020YFA0406400, No. 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts No. 11635010, No. 11735014, No. 11835012, No. 11935015, No. 11935016, No. 11935018, No. 11961141012, No. 12025502, No. 12035009, No. 12035013, No. 12061131003, No. 12192260, No. 12192261, No. 12192262, No. 12192263, No. 12192264, No. 12192265, No. 12221005, No. 12225509, No. 12235017;
the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; CAS Key Research Program of Frontier Sciences under Contracts No. QYZDJ-SSW-SLH003, No. QYZDJ-SSW-SLH040; 100 Talents Program of CAS;
Project No. ZR2022JQ02 supported by Shandong Provincial
@@
The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; European Union’s Horizon 2020 research and innovation programme under
Marie Sklodowska-Curie grant agreement under Contract No. 894790; German Research Foundation DFG under Contracts Nos. 455635585, Collaborative Research Center CRC 1044, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
Chatrchyan et al. (2013)S. Chatrchyan et al. (CMS Collaboration), “Measurement of the (3872) Production Cross Section Via Decays to in collisions at = 7 TeV,” JHEP 04, 154 (2013), arXiv:1302.3968 [hep-ex] .
Cao et al. (2020)P. Cao et al., “Design and construction of the new BESIII endcap Time-of-Flight system with MRPC Technology,” Nucl. Instrum. Meth. A 953, 163053 (2020).
Chen et al. (2000)J. C. Chen, G. S. Huang, X. R. Qi, D. H. Zhang, and Y. S. Zhu, “Event generator for and decay,” Phys. Rev. D 62, 034003 (2000).
Yang et al. (2014)R. L. Yang, R. G. Ping, and H. Chen, “Tuning and Validation of the Lundcharm Model with Decays,” Chin. Phys. Lett. 31, 061301 (2014).
Actis et al. (2010)S. Actis et al. (Working Group on Radiative Corrections, Monte Carlo Generators for Low Energies), “Quest for precision in hadronic cross sections at low energy: Monte Carlo tools vs. experimental data,” Eur. Phys. J. C 66, 585–686 (2010), arXiv:0912.0749 [hep-ph] .